OpenStudy (anonymous):

6. Name the radius and the equation of the circle with center at (3,-2), which also passes through the point . (0,2)

OpenStudy (anonymous):

the general equation is (x - h)^2 + (y - k)^2 = r^2. and the center is (h, k). so plug in for h and k into the equation, plug in the point, and solve for r. then plug in h and k and r to get the equation.

OpenStudy (anonymous):

since the radius can be find out by distance formula between centre(3,-2) and given [pont on circle(0,2)If the point is on the circle, then the distance from center to pt. is (5) the radius. The equation of the circle then becomes: (x -3)² + (x -(-2))² = 5² = = = 2) The radius can be determined after finding half of the diameter (you're given the endpts of the diameter). The diameter is 10; the radius is 5; the center is (midpt of diameter: (1, 1)). So your circle is (x -1)² + (y -1)² = 5²